# Gravitational wave probes of particle dark matter: a review

### v4

@@ III.3.5 Constraints on axions through polarization measurements
Recent analyses of European pulsar timing array data, particularly from the Parkes Pulsar Polarization Array (PPPA), have performed the first dedicated searches for these axion-induced polarization oscillations Xue et al. (2024); Porayko et al. (2025). By tracking polarization angles over long time baselines, these searches place upper limits on the axion-photon coupling constant, $g_{a\gamma}$, for axion masses in the ultra-light regime.
@@
Figure 12: Taken from Xue et al. (2024). Upper limits from analyses of European pulsar timing array data on the coupling of axion-like DM to photons. $\rho_{\text{DM}}=0.4\,\textrm{GeV}/\textrm{cm}^{3}$. The physics behind this form of DM has been discussed in Section II.2. The blue and red curves correspond to the limits derived using the auto-correlation-only and full-correlation signal models, respectively. Excluded portions of the axion mass /coupling parameter space lie above these curves, and constraints from different experiments are also shown. Key timescales are shown with vertical dotted lines, included $T_{\text{obs}}=18$ years. Smoothed constraints are shown as dashed lines in cyan, yellow, green, and purple, representing limits from polarization angle (PA) variations observed in the Crab Nebula by POLARBEAR Adachi et al. (2024), analyses of PPTA DR1 and Crab Nebula data from QUIJOTE Castillo et al. (2022), measurements of CMB polarization by SPT-3G Ferguson et al. (2022), and studies of axion-like DM-induced washout effects during recombination from the Planck mission Fedderke et al. (2019), respectively. Additional constraints from the CAST experiment Anastassopoulos et al. (2017), as well as observations of SN1987A Payez et al. (2015) and X-ray spectral distortions in quasar H1821+643 Reynés et al. (2021), are shown as black dashed lines. The lower panel shows the Bayes factors (BFs, essentially, the preference for DM to be in the data versus noise) for each of the scenarios. The BF stays around zero, although some significant spikes are apparent, which has been found to be due to imprecise noise modelling, not astrophysical signals.
@@ III.4 Prospects for detecting the coupling of ultralight DM to GW interferometers
Of course, ultralight dark matter could take on any mass a priori, and ground-based GW detectors only allow us to probe a small region of this parameter space, from $\sim 10^{-14}-10^{-12}$ eV. Space-based GW detectors will permit being sensitive to masses a few orders of magnitude smaller, while high-frequency gravitational-wave (HFGW) detectors could see DM particles a few orders of magnitude larger Aggarwal et al. (2021). However, it should be noted that when placing constraints on ultralight dark matter, the stochasticity of the signal could affect the limits: if we happen to “get unlucky” and observe the signal when the phase offset reduces the signal amplitude to zero, as can be seen at the zero points in Fig. 3, we would have no constraint. Refs. Nakatsuka et al. (2023); Boddy et al. (2025) have shown that for ground- and space-based interferometers, that this effect has an $\mathcal{O}(1)$ impact on upper limits, but needs to be mitigated for pulsar timing arrays by averaging over the signals coming from many pulsars.
The prospects for future ground-based GW interferometers, such as ET and Cosmic Explorer, to observe ultralight dark matter interactions are bright. Within a decade or so of observation, such detectors could actually be competitive at masses between $10^{-15},10^{-14}]$ eV, thus permitting an extremely strong probe of new physics Pierce et al. (2018); Morisaki et al. (2021).
@@ IV Macroscopic DM transiting through GW interferometers
As mentioned in Section I, the mass of DM could span many orders of magnitude. In particular, macroscopic dark matter, with masses in the range $M_{\rm DM}\sim[1,10^{9}]$ kg, has been proposed as a viable candidate Hall et al. (2018). These candidates could include composite objects such as nuclear-density nuggets Itoh (1970); Witten (1984) or Q-balls Coleman (1985),
@@
For $M_{\rm DM}=1$ kg, this corresponds to a flux of $\Phi\sim n_{\rm DM}v_{0}\sim 3\times 10^{-10}$ km-2 s-1, assuming a typical DM velocity $v_{0}\sim 220$ km/s. Moreover, the size of DM, $r_{\rm DM}$, would be tiny with respect to $L_{\rm DM}$: $r_{\rm DM}\sim(M_{\rm DM}/(\frac{4}{3}\pi\rho_{\rm nugget}))^{1/3}\sim 3\mu\text{m}$ for $\rho_{\rm nugget}\sim 10^{17}\text{ kg/m}^{3}$. This implies that one transient event per year would be expected in a detector with an effective cross-sectional area777the area in the plane perpendicular to the DM flux within which a passing object would produce a transient of order a few hundred km2, i.e. DM would have an impact parameter of $\sim 10$ km, which is of the same order as the length of the interferometer arms (4 km). However, gravitationally interacting DM would not be able to produce such an event rate Hall et al. (2018); Du et al. (2023), so we will need to consider interactions of DM with the standard model to enhance the event rate. In Section IV.1, we will discuss the physics of Yukawa-like interactions that macroscopic DM can have with the standard model, which will enhance the strength of the transiting DM object. Then, in Section IV.2, we show the projected constraints on Yukawa-like interactions using future data from ground-based GW interferometers.
@@ IV.1 Yukawa-like interactions
where $\delta_{i,j}$ represents the coupling strengths between particles $i,j$, and $i,j$ could be either standard-model or DM particles. $r$ is a length scale, $\lambda$ is the screening length, $s$ is the spin of the interaction mediator, and $M_{i,j}$ are the two masses. By noting that the physical size of the detector and range of the interaction force are much larger than $r_{\rm DM}$, but much smaller than the separation between DM particles $L_{\rm DM}$, the interaction can be modeled in terms of an effective geometric cross section and derived using the Born approximation Born (1926):
@@
While $|\delta_{\rm SM}|<5\times 10^{-4}$ has been strongly constrained by equivalence principle experiments Adelberger et al. (2009), the DM coupling constant could be much larger, as the only constraints come from how DM self-interaction would have influenced structure formation Spergel and Steinhardt (2000) and collisions between galaxies Harvey et al. (2015). In particular, the lack of observed deceleration of DM in the bullet cluster collision constrained the cross-section of DM to be less than $\sim 1$ cm2/g Randall et al. (2008), which provides limits on the coupling constant of self-interacting DM to be:
@@
in which $\log\left[\frac{\lambda}{r_{\text{DM}}}\right]=5$, chosen only as a benchmark, corresponding to a force range much larger $r_{\rm DM}$. Based on the current upper limits on $\delta_{\rm SM}$ and $\delta_{\rm DM}$, could allow an observation of a transit event in GW interferometers. Furthermore, $\delta_{\rm DM}\gg$ 1 would actually alleviate some problems with the cold DM scenario, including overly dense areas of dwarf galaxies found in simulations, and a prediction that the Milky Way produces more stars than are observed Boylan-Kolchin et al. (2011).
@@ IV.2 Projected constraints on Yukawa-like interactions
In order to search for transiting macroscopic DM, it has been proposed in Hall et al. (2018) to use matched filtering. To do so, firstly we must calculate the matched filtering SNR $\rho$:
@@
Here, $a(f)$ is the Fourier transform of the differential acceleration $a(t)$ between the test masses, and $S_{n}(f)$ is the power spectral density of the detector. The differential acceleration $a(t)$ arises from the DM transiting through the interferometer test masses, and is calculated with the difference in accelerations between the input mirror and end mirror Saulson (2017) in each arm. The form of $a(t)$ can be derived by taking the gradient of Eq. 81:
@@
and $x_{\rm DM}$, $y_{\rm DM}$ and $z_{\rm DM}$ are the spatial coordinates of the DM object. The acceleration depends directly on the coupling of DM to standard model particles; thus, the SNR will be proportional to this coupling as well. In searches with GW data, a threshold must be set that delineates candidates that are most likely noise with those that could be of astrophysical origin. To estimate the number of detectable transiting DM events, an SNR threshold of $\rho>8$ is applied, which is standard in GW searches Maggiore (2008), and the detection rate as a function of DM mass, and coupling $g\equiv\delta_{\rm SM}\delta_{\rm DM}$ is computed:
@@
Figure 13: Taken from Hall et al. (2018). Expected number of macroscopic DM transit events per year detectable by advanced LIGO at design sensitivity, as a function of DM mass, screening length $\lambda$, and coupling $g=\delta_{\rm SM}\delta_{\rm DM}$. The physics behind this form of DM has been discussed in Section IV.1. The blue, red and green colors indicate different screening lengths $\lambda$ of the Yukawa-like force. The three panels show the rate of transits for DM objects with three different masses. The rate is calculated using Eq. 95 and the preceding equations, assuming a detection threshold on the SNR of 8. The threshold fixes the minimum detectable coupling accessible to a matched filtering search.
Further work in Du et al. (2023) has expanded on this analysis, incorporating additional effects of DM transits, such as the Doppler effect, Shapiro delay, and Einstein delay, each of which contributes to the strain observed in the interferometer. These effects arise from the interaction of the DM object with the interferometer’s test masses, leading to small deviations in the expected strain signal.
@@
While these gravitational effects are real, they are typically very weak unless DM interacts with standard-model particles. To make the effects measurable by GW interferometers, Ref. Du et al. (2023) confirmed that DM must couple with standard-model particles, agreeeing with the findings of Hall et al. (2018). The strength of this coupling is given in Du et al. (2023) by the effective parameter $\tilde{\alpha}\sim\delta_{\rm SM}\delta_{\rm DM}$, where $\tilde{\alpha}=g$ in Hall et al. (2018). Thus, detecting these weak gravitational signals requires a sufficiently large value of $\tilde{\alpha}$, which can be constrained by observations of DM transits.
Fig. 14 shows the projected 90% sensitivity on $\tilde{\alpha}$, which can be probed by GW detectors for macroscopic DM. The results, shown for two different screening lengths ($\lambda=1\,\text{m}$ and $10^{6}\,\text{m}$), are compared to existing constraints from the Bullet Cluster Spergel and Steinhardt (2000), MICROSCOPE Bergé et al. (2018), and neutron-star kinetic heating experiments Gresham et al. (2023). While GW detectors may not surpass these existing constraints, they provide complementary information, especially if part of the DM population interacts with the standard model via a new fifth force.
Figure 14: Taken from Du et al. (2023). Projected sensitivity at 90% confidence to the standard-model/DM interaction parameter $\tilde{\alpha}\sim\delta_{\rm SM}\delta_{\rm DM}$ from macroscopic DM transits through several GW detectors. The physics behind this form of DM has been discussed in Section IV.1. The figure compares the projected sensitivity to existing constraints, including those from the Bullet Cluster, MICROSCOPE, and neutron-star kinetic heating, for two choices of screening length. Here, $T_{\text{obs}}=1$, and $\lambda=1$ m and $\lambda=10^{6}$ m, and an SNR threshold of 2 are assumed. The gray line represents combined limits from Bullet Cluster observations Spergel and Steinhardt (2000); Kahlhoefer et al. (2014) and the MICROSCOPE experiment Bergé et al. (2018); Fayet (2019). Purple and red lines indicate bounds from neutron star kinetic heating Gresham et al. (2023), with and without additional short-range DM–baryon interactions, respectively.
While deterministic signals might be challenging to detect due to the weak gravitational interactions, there is also potential for detecting a stochastic background arising from numerous transiting DM events. This stochastic background could be visible if the Yukawa interaction is sufficiently strong, as explored in Hall et al. (2018); Du et al. (2023). Future work will focus on whether a stochastic signal could be distinguished from the background noise in upcoming GW observations.
@@ V GWs from ultralight boson clouds around rotating black holes
If new ultralight bosons [^8] In the context of boson clouds around rotating black holes, this new particle could, but need not, be DM exist, quantum fluctuations could allow them to pop into existence near a rotating BH. Some bosons will fall in; others will scatter off, and the extent to which either of these scenarios happens depends on the mass and spin of both the black hole and the new boson. If the following so-called “superradiance condition”, which relates the BH angular frequency $\Omega$ to the boson angular frequency $\omega$ via the magnetic quantum number $m$, is met Brito et al. (2015):
@@
More quantitatively, solving the Klein-Gordon equation for a Kerr BH hole in the presence of a massive scalar field results in a Schrondinger-like equation that exhibits a $1/r$ potential, resulting from the gravitational interaction between the bosons and the BH. The solution of this equation shows that the energy states of the boson cloud are quantized analogously to those of the hydrogen atom, allowing us to describe the boson cloud as a “gravitational atom in the sky” Baumann et al. (2019a).
@@
The cloud will continue to build up until the BH is spun down enough such that the condition in Eq. 96 is no longer satisfied. In particular, the timescales $\tau_{\rm inst}^{\rm(s)}$ and $\tau_{\rm inst}^{\rm(v)}$ for which this occurs scalar and vector clouds, respectively, are Isi et al. (2019) [^9] Note that we do not consider tensor boson clouds in this work:
@@
This new boson could have a spin of 0, 1 or 2, where spin-0 could correspond to the pseudo-scalar QCD axion (or axion-like) particle Brito et al. (2015, 2020), spin-1 could indicate dark photons arising through kinetic mixing with the ordinary photon Siemonsen et al. (2023a) [^10] Note that this dark photon is not the dark photon described in Section II.4., or spin-2 tensor bosons that could arise from modified theories of gravity Brito et al. (2020). After the cloud builds up in an isolated BH system, GWs can be emitted in three different ways: (1) through boson-boson annihilation, (2) through boson energy-level transitions, and (3) after a “bosenova”, all of which will be discussed in Sections V.1, V.2 and V.4. Furthermore, the GW frequency is fixed by the boson mass, which could in principle take on any ultralight mass; thus, it is important to note that GW interferometers, from pulsar timing arrays to HFGW instruments, will allow an extensive mass range to be probed, from $10^{-22},10^{-9}]\,\mathrm{eV}$ Aggarwal et al. (2021); Brzeminski et al. (2025); Arvanitaki et al. (2015a); Brito et al. (2017); Ng et al. (2020).
It has been shown, under a wide range of assumptions of BH populations, spins, redshift distributions, etc., that such systems could be detectable by CW searches in advanced LIGO, Virgo and KAGRA data and/or in future detectors, such as ET and Cosmic Explorer Arvanitaki et al. (2015a); Zhu et al. (2020); and if such sources are not seen, particular boson masses can be disfavored – this will be the subject of Section VI. Furthermore, both galactic BHs Palomba et al. (2019) and remnants of mergers of BHs outside the galaxy could also be detected, the latter of which, in the scalar case, only being visible in ET and Cosmic Explorer Arvanitaki et al. (2010, 2015a); Isi et al. (2019), which will be discussed more in Section VI.1. Annihilating vector boson cloud systems, however, emit stronger, though shorter, GW signals Brito et al. (2015); Baryakhtar et al. (2017); Siemonsen et al. (2023b); Jones et al. (2023, 2025), and potentially may contain electromagnetic counterparts if the vector boson kinetically mixes with the ordinary photon Siemonsen et al. (2023a).

### v5

@@ III.3.5 Constraints on axions through polarization measurements
Recent analyses of European pulsar timing array data, particularly from the Parkes Pulsar Polarization Array (PPPA), have performed the first dedicated searches for these axion-induced polarization oscillations Xue:2024zjq; EPTA:2024gxu. By tracking polarization angles over long time baselines, these searches place upper limits on the axion-photon coupling constant, $g_{a\gamma}$, for axion masses in the ultra-light regime.
@@
Figure 12: Taken from Xue:2024zjq. Upper limits from analyses of European pulsar timing array data on the coupling of axion-like DM to photons. $\rho_{\text{DM}}=0.4\,\textrm{GeV}/\textrm{cm}^{3}$. The physics behind this form of DM has been discussed in Section II.2. The blue and red curves correspond to the limits derived using the auto-correlation-only and full-correlation signal models, respectively. Excluded portions of the axion mass /coupling parameter space lie above these curves, and constraints from different experiments are also shown. Key timescales are shown with vertical dotted lines, included $T_{\text{obs}}=18$ years. Smoothed constraints are shown as dashed lines in cyan, yellow, green, and purple, representing limits from polarization angle (PA) variations observed in the Crab Nebula by POLARBEAR POLARBEAR:2024vel, analyses of PPTA DR1 and Crab Nebula data from QUIJOTE Castillo:2022zfl, measurements of CMB polarization by SPT-3G SPT-3G:2022ods, and studies of axion-like DM-induced washout effects during recombination from the Planck mission Fedderke:2019ajk, respectively. Additional constraints from the CAST experiment CAST:2017uph, as well as observations of SN1987A Payez:2014xsa, updated in Fiorillo:2025gnd, and X-ray spectral distortions in quasar H1821+643 Reynes:2021bpe, are shown as black dashed lines. The lower panel shows the Bayes factors (BFs, essentially, the preference for DM to be in the data versus noise) for each of the scenarios. The BF stays around zero, although some significant spikes are apparent, which has been found to be due to imprecise noise modelling, not astrophysical signals.
@@ III.4 Prospects for detecting the coupling of ultralight DM to GW interferometers
Of course, ultralight dark matter could take on any mass a priori, and ground-based GW detectors only allow us to probe a small region of this parameter space, from $\sim 10^{-14}-10^{-12}$ eV. Space-based GW detectors will permit being sensitive to masses a few orders of magnitude smaller, while high-frequency gravitational-wave (HFGW) detectors could see DM particles a few orders of magnitude larger Aggarwal:2020olq. However, it should be noted that when placing constraints on ultralight dark matter, the stochasticity of the signal could affect the limits: if we happen to “get unlucky” and observe the signal when the phase offset reduces the signal amplitude to zero, as can be seen at the zero points in Fig. 3, we would have no constraint. Refs. Nakatsuka:2022gaf; Boddy:2025oxn have shown that for ground- and space-based interferometers, that this effect has an $\mathcal{O}(1)$ impact on upper limits, but needs to be mitigated for pulsar timing arrays by averaging over the signals coming from many pulsars.
The prospects for future ground-based GW interferometers, such as ET and Cosmic Explorer, to observe ultralight dark matter interactions are bright. Within a decade or so of observation, such detectors could actually be competitive at masses between $10^{-15},10^{-14}]$ eV, thus permitting an extremely strong probe of new physics Pierce:2018xmy; Morisaki:2020gui.
@@ IV Macroscopic DM transiting through GW interferometers
As mentioned in Section I, the mass of DM could span many orders of magnitude. In particular, macroscopic dark matter, with masses in the range $M_{\rm DM}\sim[1,10^{9}]$ kg, has been proposed as a viable candidate Hall:2016usm. These candidates could include composite objects such as nuclear-density nuggets Itoh:1970uw; Witten:1984rs or Q-balls Coleman:1985ki,
@@
For $M_{\rm DM}=1$ kg, this corresponds to a flux of $\Phi\sim n_{\rm DM}v_{0}\sim 3\times 10^{-10}$ km-2 s-1, assuming a typical DM velocity $v_{0}\sim 220$ km/s. Moreover, the size of DM, $r_{\rm DM}$, would be tiny with respect to $L_{\rm DM}$: $r_{\rm DM}\sim(M_{\rm DM}/(\frac{4}{3}\pi\rho_{\rm nugget}))^{1/3}\sim 3\mu\text{m}$ for $\rho_{\rm nugget}\sim 10^{17}\text{ kg/m}^{3}$. This implies that one transient event per year would be expected in a detector with an effective cross-sectional area777the area in the plane perpendicular to the DM flux within which a passing object would produce a transient of order a few hundred km2, i.e. DM would have an impact parameter of $\sim 10$ km, which is of the same order as the length of the interferometer arms (4 km). However, gravitationally interacting DM would not be able to produce such an event rate Hall:2016usm; Du:2023dhk, so we will need to consider interactions of DM with the standard model to enhance the event rate. In Section IV.1, we will discuss the physics of Yukawa-like interactions that macroscopic DM can have with the standard model, which will enhance the strength of the transiting DM object. Then, in Section IV.2, we show the projected constraints on Yukawa-like interactions using future data from ground-based GW interferometers.
@@ IV.1 Yukawa-like interactions
where $\delta_{i,j}$ represents the coupling strengths between particles $i,j$, and $i,j$ could be either standard-model or DM particles. $r$ is a length scale, $\lambda$ is the screening length, $s$ is the spin of the interaction mediator, and $M_{i,j}$ are the two masses. By noting that the physical size of the detector and range of the interaction force are much larger than $r_{\rm DM}$, but much smaller than the separation between DM particles $L_{\rm DM}$, the interaction can be modeled in terms of an effective geometric cross section and derived using the Born approximation Born:1926oue:
@@
While $|\delta_{\rm SM}|<5\times 10^{-4}$ has been strongly constrained by equivalence principle experiments Adelberger:2009zz, the DM coupling constant could be much larger, as the only constraints come from how DM self-interaction would have influenced structure formation Spergel:1999mh and collisions between galaxies Harvey:2015hha. In particular, the lack of observed deceleration of DM in the bullet cluster collision constrained the cross-section of DM to be less than $\sim 1$ cm2/g Randall:2008ppe, which provides limits on the coupling constant of self-interacting DM to be:
@@
in which $\log\left[\frac{\lambda}{r_{\text{DM}}}\right]=5$, chosen only as a benchmark, corresponding to a force range much larger $r_{\rm DM}$. Based on the current upper limits on $\delta_{\rm SM}$ and $\delta_{\rm DM}$, could allow an observation of a transit event in GW interferometers. Furthermore, $\delta_{\rm DM}\gg$ 1 would actually alleviate some problems with the cold DM scenario, including overly dense areas of dwarf galaxies found in simulations, and a prediction that the Milky Way produces more stars than are observed Boylan-Kolchin:2011qkt.
@@ IV.2 Projected constraints on Yukawa-like interactions
In order to search for transiting macroscopic DM, it has been proposed in Hall:2016usm to use matched filtering. To do so, firstly we must calculate the matched filtering SNR $\rho$:
@@
Here, $a(f)$ is the Fourier transform of the differential acceleration $a(t)$ between the test masses, and $S_{n}(f)$ is the power spectral density of the detector. The differential acceleration $a(t)$ arises from the DM transiting through the interferometer test masses, and is calculated with the difference in accelerations between the input mirror and end mirror Saulson:2017jlf in each arm. The form of $a(t)$ can be derived by taking the gradient of Eq. 81:
@@
and $x_{\rm DM}$, $y_{\rm DM}$ and $z_{\rm DM}$ are the spatial coordinates of the DM object. The acceleration depends directly on the coupling of DM to standard model particles; thus, the SNR will be proportional to this coupling as well. In searches with GW data, a threshold must be set that delineates candidates that are most likely noise with those that could be of astrophysical origin. To estimate the number of detectable transiting DM events, an SNR threshold of $\rho>8$ is applied, which is standard in GW searches maggiore2008gravitational, and the detection rate as a function of DM mass, and coupling $g\equiv\delta_{\rm SM}\delta_{\rm DM}$ is computed:
@@
Figure 13: Taken from Hall:2016usm. Expected number of macroscopic DM transit events per year detectable by advanced LIGO at design sensitivity, as a function of DM mass, screening length $\lambda$, and coupling $g=\delta_{\rm SM}\delta_{\rm DM}$. The physics behind this form of DM has been discussed in Section IV.1. The blue, red and green colors indicate different screening lengths $\lambda$ of the Yukawa-like force. The three panels show the rate of transits for DM objects with three different masses. The rate is calculated using Eq. 95 and the preceding equations, assuming a detection threshold on the SNR of 8. The threshold fixes the minimum detectable coupling accessible to a matched filtering search.
Further work in Du:2023dhk has expanded on this analysis, incorporating additional effects of DM transits, such as the Doppler effect, Shapiro delay, and Einstein delay, each of which contributes to the strain observed in the interferometer. These effects arise from the interaction of the DM object with the interferometer’s test masses, leading to small deviations in the expected strain signal.
@@
While these gravitational effects are real, they are typically very weak unless DM interacts with standard-model particles. To make the effects measurable by GW interferometers, Ref. Du:2023dhk confirmed that DM must couple with standard-model particles, agreeeing with the findings of Hall:2016usm. The strength of this coupling is given in Du:2023dhk by the effective parameter $\tilde{\alpha}\sim\delta_{\rm SM}\delta_{\rm DM}$, where $\tilde{\alpha}=g$ in Hall:2016usm. Thus, detecting these weak gravitational signals requires a sufficiently large value of $\tilde{\alpha}$, which can be constrained by observations of DM transits.
Fig. 14 shows the projected 90% sensitivity on $\tilde{\alpha}$, which can be probed by GW detectors for macroscopic DM. The results, shown for two different screening lengths ($\lambda=1\,\text{m}$ and $10^{6}\,\text{m}$), are compared to existing constraints from the Bullet Cluster Spergel:1999mh, MICROSCOPE Berge:2017ovy, and neutron-star kinetic heating experiments Gresham:2022biw. While GW detectors may not surpass these existing constraints, they provide complementary information, especially if part of the DM population interacts with the standard model via a new fifth force.
Figure 14: Taken from Du:2023dhk. Projected sensitivity at 90% confidence to the standard-model/DM interaction parameter $\tilde{\alpha}\sim\delta_{\rm SM}\delta_{\rm DM}$ from macroscopic DM transits through several GW detectors. The physics behind this form of DM has been discussed in Section IV.1. The figure compares the projected sensitivity to existing constraints, including those from the Bullet Cluster, MICROSCOPE, and neutron-star kinetic heating, for two choices of screening length. Here, $T_{\text{obs}}=1$, and $\lambda=1$ m and $\lambda=10^{6}$ m, and an SNR threshold of 2 are assumed. The gray line represents combined limits from Bullet Cluster observations Spergel:1999mh; Kahlhoefer:2013dca and the MICROSCOPE experiment Berge:2017ovy; Fayet:2018cjy. Purple and red lines indicate bounds from neutron star kinetic heating Gresham:2022biw, with and without additional short-range DM–baryon interactions, respectively.
While deterministic signals might be challenging to detect due to the weak gravitational interactions, there is also potential for detecting a stochastic background arising from numerous transiting DM events. This stochastic background could be visible if the Yukawa interaction is sufficiently strong, as explored in Hall:2016usm; Du:2023dhk. Future work will focus on whether a stochastic signal could be distinguished from the background noise in upcoming GW observations.
@@ V GWs from ultralight boson clouds around rotating black holes
If new ultralight bosons [^8] In the context of boson clouds around rotating black holes, this new particle could, but need not, be DM exist, quantum fluctuations could allow them to pop into existence near a rotating BH. Some bosons will fall in; others will scatter off, and the extent to which either of these scenarios happens depends on the mass and spin of both the black hole and the new boson. If the following so-called “superradiance condition”, which relates the BH angular frequency $\Omega$ to the boson angular frequency $\omega$ via the magnetic quantum number $m$, is met Brito:2015oca:
@@
More quantitatively, solving the Klein-Gordon equation for a Kerr BH hole in the presence of a massive scalar field results in a Schrondinger-like equation that exhibits a $1/r$ potential, resulting from the gravitational interaction between the bosons and the BH. The solution of this equation shows that the energy states of the boson cloud are quantized analogously to those of the hydrogen atom, allowing us to describe the boson cloud as a “gravitational atom in the sky” Baumann:2018vus.
@@
The cloud will continue to build up until the BH is spun down enough such that the condition in Eq. 96 is no longer satisfied. In particular, the timescales $\tau_{\rm inst}^{\rm(s)}$ and $\tau_{\rm inst}^{\rm(v)}$ for which this occurs scalar and vector clouds, respectively, are Isi:2018pzk [^9] Note that we do not consider tensor boson clouds in this work:
@@
This new boson could have a spin of 0, 1 or 2, where spin-0 could correspond to the pseudo-scalar QCD axion (or axion-like) particle Brito:2015oca; Brito:2020lup, spin-1 could indicate dark photons arising through kinetic mixing with the ordinary photon Siemonsen:2022ivj [^10] Note that this dark photon is not the dark photon described in Section II.4., or spin-2 tensor bosons that could arise from modified theories of gravity Brito:2020lup. After the cloud builds up in an isolated BH system, GWs can be emitted in three different ways: (1) through boson-boson annihilation, (2) through boson energy-level transitions, and (3) after a “bosenova”, all of which will be discussed in Sections V.1, V.2 and V.4. Furthermore, the GW frequency is fixed by the boson mass, which could in principle take on any ultralight mass; thus, it is important to note that GW interferometers, from pulsar timing arrays to HFGW instruments, will allow an extensive mass range to be probed, from $10^{-22},10^{-9}]\,\mathrm{eV}$ Aggarwal:2020olq; Brzeminski:2024drp; Arvanitaki:2014wva; Brito:2017zvb; Ng:2020jqd.
It has been shown, under a wide range of assumptions of BH populations, spins, redshift distributions, etc., that such systems could be detectable by CW searches in advanced LIGO, Virgo and KAGRA data and/or in future detectors, such as ET and Cosmic Explorer Arvanitaki:2014wva; Zhu:2020tht; and if such sources are not seen, particular boson masses can be disfavored – this will be the subject of Section VI. Furthermore, both galactic BHs Palomba:2019vxe and remnants of mergers of BHs outside the galaxy could also be detected, the latter of which, in the scalar case, only being visible in ET and Cosmic Explorer Arvanitaki:2009fg; Arvanitaki:2014wva; Isi:2018pzk, which will be discussed more in Section VI.1. Annihilating vector boson cloud systems, however, emit stronger, though shorter, GW signals Brito:2015oca; Baryakhtar:2017ngi; Siemonsen:2022yyf; Jones:2023fzz; Jones:2024fpg, and potentially may contain electromagnetic counterparts if the vector boson kinetically mixes with the ordinary photon Siemonsen:2022ivj.